# Equal contribution

### v1

@@ 4.2 Architecture
It has been demonstrated in [lugmayr2020srflow] that Conditional NFs are capable of learning distributions on the ambient space that are constrained on measurement. To achieve an architecture with invertible transformations, we extend RealNVP [dinh2016density] architecture to the conditional settings by borrowing insight from [sun2021deep]. In its plain form, RealNVP architecture mainly consists of Flow steps, each containing two Affine Coupling layers. In each affine coupling layer, input signal ${\bm{x}}$ is split into two parts: ${\bm{x}}_{a}$ which stays unchanged, and ${\bm{x}}_{b}$ which is fed into the neural network. In order to invoke the condition, we simply concatenate the conditioning input ${\bm{y}}$ to the ${\bm{x}}_{b}$ as these layers serve as the main and basic building blocks of entire invertible architecture. This seemingly simple integration led to very promising results in both Euclidean and non-Euclidean geometries as will be depicted in Section 5.
@@ 4.3 Manifold
Manifolds exhibit more complex or curved structures compared to flat Euclidean spaces. In many real-world scenarios such as in environmental science, data naturally resides on these non-Euclidean geometries, and removing artifacts or denoising over these shapes becomes both necessary and hard. Traditional DIS methods usually rely on spatial domain and requires for off-the-shelf projection methods to get plausible results. To mitigate this, our work extends the application of conditional NFs to distributions lying on non-Euclidean manifolds and solve intricate inverse problems right on the surface without the need for additional renderings. By leveraging the expressive power of conditional NFs, we aim to capture the intrinsic geometry and structure of manifold data while enabling efficient inference and sampling. We treat data as a point cloud of size $V\times C$, in which $V$ and $C$ denotes the number of vertices on the discretization of mesh and dimension of signal features (e.g. $3$ in case of rgb), respectively. Several complex geometries and practical data have been considered and CNF is optimized to perform noisy inpainting and imputation tasks. Findings suggest that our framework can capture the distribution even in severe masking levels across different geometries (see Section 5).
@@ 5 Experiments
We validate our approach through various experiments, including (i) Denoising, Super Resolution, and Deblurring with CelebA face image data [liu2015faceattributes]; (ii) Inpainting on Stanford Bunny Manifold with MNIST data; and (iii) Imputation on Sphere with ERA5 [hersbach2020era5] temperature data. (i) Denoising is performed on the Euclidean in the image domain. In contrast, noisy (ii) inpainting and (iii) imputation are solved directly on the bunny and sphere manifolds. Throughout all the experiments, we use $24$ flow steps and we set the batch size to the $64$. We conduct all the training and optimization experiments on a single RTX3090 GPU instance. Our code is implemented in the JAX framework [jax2018github].
5.1 Experimental Settings
Inverse Problems on CelebA. We follow the usual formulation and adapt $32\times 32$ resolution of facial images. Data is normalized into [0, 1] range and measurement is acquired by the appropriate choices of forward operator depending on the task (See Appendix A for details). $G_{\phi}$ is optimized over the $19,962$ test images by using the forward operator and prior from diffusion models. We optimize CNF for $1$M iterations (convergency was observed earlier but continued for potential refinement).
Inpainting on Bunny MNIST. In order to demonstrate the geometric awareness of our model, we conduct experiment on Stanford Bunny Manifold. We choose the mesh resolution of $1889$ vertices and then project the [0, 1] normalized MNIST digits onto the bunny manifold [turk1994zippered]. In order to ensure the dimensionality compatibility for the models, we use $1888$ vertices and zero mask the last vertex throughout the experiments. We obtain the measurement by occluding $30\%$ of vertices randomly and adding some Gaussian noise, i.e. ${\mathcal{A}}$ is the random masking operator and $\sigma_{y}=0.1$ in (1). Our proposed CNF is trained on the test chunk of $10,000$ digit examples for $1$M iterations.
Imputation on ERA5. To show the essence and practical importance of our pipeline, we further conduct experiments on ERA5 temperature dataset. Even though data is available in a rectangular format, due to the spherical shape of Earth, it inherits some geometric information. We use $4^{\circ}$ resolution dataset with $4140$ vertices borrowed from [dupontetal5gasp] with only temperature channel as it is quite popular to analyze in the domain of generative AI. Again, due to the dimensionality, we add $20$ more vertices with a signal value of zero, and the data is [0, 1] normalized. In contrast to Bunny MNIST, we use more severe occlusion of $60\%$ random masking with additional Gaussian noise of $\sigma_{y}=0.05$. We perform the optimization of our CNF model on the test part of the dataset with $2420$ examples for $315$k iterations.
Score Networks. For all the diffusion priors VPSDE formulation has been adapted. In the case of image domain CelebA, we borrow the same score checkpoint used in the [feng2023score, feng2023efficient], which uses NCSN++ [song2021maximum] architecture and has been trained for $1$M iterations. For Bunny MNIST, we adapt the $1$D formulation of DDPM [ho2020denoising] and train the score network for $500$k, at which the convergence was clearly observed through the generated samples. In the case of spherical weather data, we followed the same strategy as Bunny MNIST but achieved convergence of score network earlier at $360$k iterations.
@@ 5.1 Experimental Settings
MCG [chung2022improving] - - - 19.85 26.69 0.0024 16.16 27.52 0.871
Noise2Score [kim2021noise2score] 0.0172 24.36 0.871 - - - - - -
DPS [chung2023diffusion] 16.95 27.93 0.932 19.39 28.03 0.0017 15.36 28.95 0.953
Proposed ($N=1$) 0.0021 23.37 0.836 0.0021 24.04 0.0041 0.0012 33.17 0.883
Proposed ($N=128$) 0.0035 25.82 0.901 0.0035 24.88 0.0034 0.0018 34.61 0.959
@@ 5.2 Results
In this section, we provide the general results of each different task described above. We compare our model with the various baselines including, DPS [chung2023diffusion], MCG [chung2022improving], and Noise2Score [kim2021noise2score]. It should be noted that MCG and DPS are identical for denoising, and Noise2Score is only applicable to denoising. In such cases, we do not report the metrics. We also demonstrate the comparisons and results against Feng et al. [feng2023efficient]. Finally, we experimentally confirm the robustness of our model across different unseen data or datasets. For evaluation purposes, we use peak signal-to-noise ratio (PSNR) and structural-similarity-index-measure (SSIM) which are widely used to assess the performance of inverse solvers with the ground truth and reconstructed signals. We further evaluate Fréchet Inception Distance (FID) to showcase the perceptual quality of generated samples. As the proposed method can sample multiple different posterior samples with a single forward pass, and this process is easily parallelizable, we report two different types for the proposed method. One by taking a single posterior sample ($N=1$), and another by taking 128 posterior samples and taking the mean ($N=128$).
5.2.1 General Results.
[…]
5.2.2 Comparing with DIS and Noise2Score.
Both DPS [chung2023diffusion] and MCG [chung2022improving] leverage the pre-trained diffusion model to sample from the posterior distribution. However, these methods require thousands of NFEs to achieve stable performance. The required time for DPS and MCG is reported in Tab. 2, 3. When decreasing the NFE as shown in Fig. 3 (a), PSNR heavily degrades and eventually diverges when we take an NFE value of less than 30. Our method achieves competitive performance even with a single NFE. Moreover, it is shown in Fig. 3 (b) that even slightly incorrectly choosing the step size parameter leads to a large degradation in performance, whereas our method is free from such cumbersome hyperparameter tuning. Finally, it is shown in Fig. 3 (c) that DPS collapses to the mean of the prior distribution, altering the content of the measurement heavily when we take a smaller number of NFEs.
Figure 3: Comparison of our method against DPS [chung2023diffusion] on celebA denoising. (a) NFE vs. PSNR plot, (b) step size (used in DPS only) vs. PSNR plot, (c) representative results by varying the NFE.
It is worth mentioning that Noise2Score [kim2021noise2score] is applicable for one-step denoising of the measurements by leveraging the Tweedie’s formula. However, as discussed in Sec. 6, our method is generally applicable to a wide class of inverse problems, whereas the applicability of Noise2Score is limited.
@@
Proposed 26.34 0.923 56.11 23.81 0.860 86.23 0.0035
MCG [chung2022improving] 27.38 0.928 75.63 22.64 0.859 100.3 17.27
DPS [chung2023diffusion] 27.40 0.928 74.84 24.44 0.865 89.92 16.95
@@
5.2.3 Comparing with score prior method.
Compared to the exact score prior [feng2023score], surrogate counterpart [feng2023efficient] presents $100$ times faster approach along with competitive or slightly better results in terms of quality. Despite being fast in terms of optimization of NF, [feng2023efficient] still requires training the network for a considerable amount of time for every single measurement. We observed that under same conditions, conditional NF does not increase the complexity and training stage takes 0.15 seconds which is 0.14 seconds in case of unconditional version. We further sample a random point from test data of celebA and optimize unconditional NF with the same configurations as ours on this single measurement. NF trained solely on this data reaches 23.75dB in PSNR score, which is almost same as our result of 23.43dB on this measurement. All these confirms that under same conditions, our approach of introducing conditions can simply achieve best results being also amortized for plenty of measurements.
5.2.4 Generalizability across datasets and Blind Inverse Problems.
We further observe that our optimized framework can be used on unseen data as well. Table 4 and Figure 7 depicts that we achieve similar quantitative and qualitative reconstruction results when we sample from unseen celebA or ERA5 validation datasets. Note that score network is trained on train signals and CNF optimization has been conducted on test signals, i.e. validation is totally hidden to both teacher and student models. More strongly, our approach can also be leveraged on various datasets as pre-optimized inverse solver. To this end, we use our celebA optimized CNF model to perform Denoising task on the FFHQ [karras2019style] dataset. Same Table and Figure show that our model can remove the noise artifacts with a similar performance as it does on original data, confirming the generalizibility feature.
We further observed that our amortized inference approach can work in the absence of forward operator. In other words, we can perform blind inverse problems through our amortized posterior sampling. We used various imputation levels between 30% to 60% for ERA5 dataset, and conducted experiments with random choice of imputation in a blind manner. As a result, Fig. 5shows that results as good as the original inverse solver with the known forward operation (all levels with more than 33 PSNR value).
@@
[…]
Table 4: Our approach is robust against unseen data samples and even generalizable accross different datasets once it is optimized. For the first $2$ columns, we use validation part of datasets and feed-forward our CNF on this totally unseen data. For the third column, we even show that our pre-optimized CNF can be leveraged to restore back the noised data samples from across various datasets. All the quantitative results align with their counterparts in Table 2 that confirms robustness and generalization ability of our pipeline which was not possible before.
Figure 6: Figure demonstrating the visual results of robustness and generalization ability of our approach. First and second rows show the results on unseen validation data, and third row depicts generalization to another dataset. Corresponding quantitative analysis can be found in Table 4.
@@ 6 Discussion
We show a first proof of concept that we can construct a one-step posterior sampler that generalizes across any measurements in an unsupervised fashion (only having access to the measurements ${\bm{y}}$). Notably, our method extends to wide use cases with minimal constraints: 1) the operator ${\mathcal{A}}$ can be arbitrarily complex and non-linear, as in DPS [chung2023diffusion], unlike many recent DIS that requires linearity of the operator [wang2023zeroshot, chung2024decomposed, zhu2023denoising]; 2) Training of the sampler can be done without any strict conditions on the measurement, unlike recent unsupervised score training methods that require i.i.d. measurement conditions with the same randomized forward operator [daras2023ambient, kawar2023gsure]. We opted for simplicity in the architecture design of $G_{\phi}$, and avoided introducing inductive bias of spatial information by taking a vectorized input, potentially explaining the slight background noise in the reconstructions. Further optimization in the choice of network architecture is left as a future direction of study.
@@ 7 Conclusion
In this work, we propose to use a conditional NF for a VI-based optimization strategy to train a one-step posterior sampler, which implicitly samples from the posterior distribution defined from the pre-trained diffusion prior. We show that our method is highly generalizable, being able to reconstruct samples that are not seen during training, applicable to diverse forward measurements, and types of data, encompassing standard Euclidean geometry as well as data on general Riemannian manifolds. We believe that our work can act as a cornerstone for developing a fast, practical posterior sampler that distills the diffusion prior.
A Appendix
@@ A.1 Reproducibility and Details of Parameters
We further provide all the necessary details to replicate the results with our proposed method. Table 5 demonstrates the details of tasks and datasets along with the parameter choices for both prior score network training and Conditional NF optimization. Note that, we have validated our approach through $3$ different inverse problems on the image dataset (celebA), where noise level was set to 0.1 for denoising and 0.01 for Super-Resolution and Gaussian Deblurring. In addition to this, we also submit all the codes as a supplementary material to reproduce the results shown throughout the paper.
@@
(C) #optimization iters 925k 1M 315k
Table 5: Different configurations of hyperparameter choices for varying datasets and manifolds learned by our model. (S) and (C) denotes the parameter choices for score network and CNF optimization, respectively.
A.2 Qualitative comparisons with baselines

### v2

@@ 4.2 Architecture
It has been demonstrated in (lugmayr2020srflow) that Conditional NFs are capable of learning distributions on the ambient space that are constrained on measurement. To achieve an architecture with invertible transformations, we extend RealNVP (dinh2016density) architecture to the conditional settings by borrowing insight from (sun2021deep). In its plain form, RealNVP architecture mainly consists of Flow steps, each containing two Affine Coupling layers. In each affine coupling layer, input signal ${\bm{x}}$ is split into two parts: ${\bm{x}}_{a}$ which stays unchanged, and ${\bm{x}}_{b}$ which is fed into the neural network. In order to invoke the condition, we simply concatenate the conditioning input ${\bm{y}}$ to the ${\bm{x}}_{b}$ as these layers serve as the main and basic building blocks of entire invertible architecture. This seemingly simple integration led to very promising results in both Euclidean and non-Euclidean geometries as will be depicted in Section 5.
@@ 4.3 Manifold
Many real-world datasets, particularly in environmental science, naturally reside on non-Euclidean geometries, making inverse problems challenging. Our work extends conditional normalizing flows (CNFs) to distributions on non-Euclidean manifolds, enabling direct solving of inverse problems on these surfaces without additional rendering steps. We represent manifold data as point clouds of size $V\times C$, where $V$ is the number of vertices in the mesh discretization and $C$ is the dimension of signal features. By leveraging the expressive power of CNFs, our approach captures the intrinsic geometry and structure of manifold data while enabling efficient inference and sampling. Our framework can handle complex geometries and severe masking levels across different manifolds, as demonstrated in our experiments with noisy inpainting and imputation tasks (see Section 5).
@@ 5 Experiments
We validate our approach through various experiments, including (i) Denoising, Super Resolution (SR), and Deblurring with CelebA face image data (liu2015faceattributes); (ii) Inpainting on Stanford Bunny Manifold with MNIST data; and (iii) Imputation on Sphere with ERA5 (hersbach2020era5) temperature data. (i) Denoising, SR, and Deblurring are performed on the Euclidean in the image domain. In contrast, noisy (ii) inpainting and (iii) imputation are solved directly on the bunny and sphere manifolds. Throughout all the experiments, we use $24$ flow steps and we set the batch size to the $64$. We conduct all the training and optimization experiments on a single RTX3090 GPU instance. Our code is implemented in the JAX framework (jax2018github). Necessary details for the experiment settings are given in the Appendix A.1 & A.2.
@@
MCG (chung2022improving) - - - 19.85 26.69 0.0024 16.16 27.52 0.871
Noise2Score (kim2021noise2score) 0.0172 24.36 0.871 - - - - - -
DPS (chung2023diffusion) 16.95 27.93 0.932 19.39 28.03 0.0017 15.36 28.95 0.953
APS (ours) ($N=1$) 0.0021 23.37 0.836 0.0021 25.97 0.0032 0.0012 33.17 0.883
APS (ours) ($N=128$) 0.0035 25.82 0.901 0.0035 26.72 0.0022 0.0018 34.61 0.959
@@ 5.1 Results
In this section, we provide the general results of each different task described above. We compare APS with the various baselines including, DPS (chung2023diffusion), MCG (chung2022improving), and Noise2Score (kim2021noise2score) (see Appendix A.4.1). It should be noted that MCG and DPS are identical for denoising, and Noise2Score is only applicable to denoising. In such cases, we do not report the metrics. We also demonstrate the comparisons and results against Feng et al. (feng2024variational). Finally, we experimentally confirm the robustness of APS across different unseen data or datasets (see Appendix A.4.2). For evaluation purposes, we use peak signal-to-noise ratio (PSNR) and structural-similarity-index-measure (SSIM) which are widely used to assess the performance of inverse solvers with the ground truth and reconstructed signals. We further evaluate Fréchet Inception Distance (FID) to showcase the perceptual quality of generated samples. As the proposed method, APS, can sample multiple different posterior samples with a single forward pass, and this process is easily parallelizable, we report two different types for the proposed method. One by taking a single posterior sample ($N=1$), and another by taking 128 posterior samples and taking the mean ($N=128$).
@@
MCG (chung2022improving) 27.38 0.928 75.63 22.64 0.859 100.3 17.27
DPS (chung2023diffusion) 27.40 0.928 74.84 24.44 0.865 89.92 16.95
APS (ours) 26.34 0.923 56.11 23.81 0.860 86.23 0.0035
@@
5.1.1 General Results
@@
5.1.2 Comparing with score prior method
Compared to the exact score prior (feng2023score), surrogate counterpart (feng2024variational) presents $100$ times faster approach along with competitive or slightly better results in terms of quality. Despite being fast in terms of optimization of NF, (feng2024variational) still requires training the network for a considerable amount of time for every single measurement. We observed that under same conditions, conditional NF does not increase the complexity and training stage takes 0.15 seconds which is 0.14 seconds in case of unconditional version. We further sample a random point from test data of celebA and optimize unconditional NF with the same configurations as ours on this single measurement. NF trained solely on this data reaches 23.75dB in PSNR score, which is almost same as our result of 23.43dB on this measurement. All these confirms that under same conditions, APS can simply achieve best results being also amortized for plenty of measurements.
@@ 6 Discussion
We show a first proof of concept that we can construct a one-step posterior sampler that generalizes across any measurements in an unsupervised fashion (only having access to the measurements ${\bm{y}}$). Notably, APS extends to wide use cases with minimal constraints: 1) the operator ${\mathcal{A}}$ can be arbitrarily complex and non-linear, as in DPS (chung2023diffusion), unlike many recent DIS that requires linearity of the operator (wang2023zeroshot; chung2024decomposed; zhu2023denoising); 2) training of the sampler can be done without any strict conditions on the measurement, unlike recent unsupervised score training methods that require i.i.d. measurement conditions with the same randomized forward operator (daras2023ambient; kawar2023gsure); 3) method can be generalized into different geometries and datasets in a blind manner, unlike recent DIS methods which require to know forward operator during sampling (chung2023solving; mardani2023variational). We opted for simplicity in the architecture design of $G_{\phi}$, and avoided introducing inductive bias of spatial information by taking a vectorized input, potentially explaining the slight background noise in the reconstructions. Further optimization in the choice of network architecture is left as a future direction of study.
@@ 7 Conclusion
In this work, we propose to use a conditional NF for a VI-based optimization strategy to train a one-step posterior sampler, which implicitly samples from the posterior distribution defined from the pre-trained diffusion prior. We show that APS is highly generalizable, being able to reconstruct samples that are not seen during training, applicable to diverse forward measurements, and types of data, encompassing standard Euclidean geometry as well as data on general Riemannian manifolds. We believe that our work can act as a cornerstone for developing a fast, practical posterior sampler that distills the diffusion prior.
Appendix A Appendix
@@ A.1 Reproducibility and Details of Parameters
We further provide all the necessary details to replicate the results with our proposed method. Tab. 4 demonstrates the details of tasks and datasets along with the parameter choices for both prior score network training and Conditional NF optimization. Note that, we have validated our approach through $3$ different inverse problems on the image dataset (celebA), where noise level was set to 0.1 for denoising and 0.01 for Super-Resolution and Gaussian Deblurring. In addition to this, we also submit all the codes as a supplementary material to reproduce the results shown throughout the paper.
@@
(C) #optimization iters 1M 1.5M 315k
Table 4: Different configurations of hyperparameter choices for varying datasets and manifolds learned by APS. (S) and (C) denotes the parameter choices for score network and CNF optimization, respectively.
A.2 Experimental Settings
Inverse Problems on CelebA. We follow the usual formulation and adapt $32\times 32$ resolution of facial images. Data is normalized into [0, 1] range and measurement is acquired by the appropriate choices of forward operator depending on the task (See Appendix A for details). $G_{\phi}$ is optimized over the $19,962$ test images by using the forward operator and prior from diffusion models. We optimize APS for $1$M iterations for all different tasks (convergency was observed earlier but continued for potential refinement).
Inpainting on Bunny MNIST. In order to demonstrate the geometric awareness of our model, we conduct experiment on Stanford Bunny Manifold. We choose the mesh resolution of $1889$ vertices and then project the [0, 1] normalized MNIST digits onto the bunny manifold (turk1994zippered). In order to ensure the dimensionality compatibility for the models, we use $1888$ vertices and zero mask the last vertex throughout the experiments. We obtain the measurement by occluding $30\%$ of vertices randomly and adding some Gaussian noise, i.e. ${\mathcal{A}}$ is the random masking operator and $\sigma_{y}=0.1$ in (1). APS is optimized on the test chunk of $10,000$ digit examples for $1.5$M iterations.
Imputation on ERA5. To show the essence and practical importance of our pipeline, we further conduct experiments on ERA5 temperature dataset. Even though data is available in a rectangular format, due to the spherical shape of Earth, it inherits some geometric information. We use $4^{\circ}$ resolution dataset with $4140$ vertices borrowed from (dupontetal5gasp) with only temperature channel as it is quite popular to analyze in the domain of generative AI (dupont2021gasp; dupont2022functa). Again, due to the dimensionality, we add $20$ more vertices with a signal value of zero, and the data is [0, 1] normalized. In contrast to Bunny MNIST, we use more severe occlusion of $60\%$ random masking with additional Gaussian noise of $\sigma_{y}=0.05$. We perform the optimization of APS on the test part of the dataset with $2420$ examples for $315$k iterations.
Score Networks. For all the diffusion priors, VPSDE formulation has been adapted. In the case of image domain CelebA, we borrow the same score checkpoint used in the (feng2023score; feng2024variational), which uses NCSN++ (song2021maximum) architecture and has been trained for $1$M iterations. For Bunny MNIST, we adapt the $1$D formulation of DDPM (ho2020denoising) and train the score network for $500$k, at which the convergence was clearly observed through the generated samples. In the case of spherical weather data, we followed the same strategy as Bunny MNIST but achieved convergence of score network earlier at $360$k iterations.
A.3 Qualitative comparisons with baselines
@@
A.4 Further Experimental Results
A.4.1 Comparing with DIS and Noise2Score
Both DPS (chung2023diffusion) and MCG (chung2022improving) leverage the pre-trained diffusion model to sample from the posterior distribution. However, these methods require thousands of NFEs to achieve stable performance. The required time for DPS and MCG is reported in Tab. 2, 3. When decreasing the NFE as shown in Fig. 5 (a), PSNR heavily degrades and eventually diverges when we take an NFE value of less than 30. APS achieves competitive performance even with a single NFE. Moreover, it is shown in Fig. 5 (b) that even slightly incorrectly choosing the step size parameter leads to a large degradation in performance, whereas our method is free from such cumbersome hyperparameter tuning. Finally, it is shown in Fig. 5 (c) that DPS collapses to the mean of the prior distribution, altering the content of the measurement heavily when we take a smaller number of NFEs.
Figure 5: Comparison of our method against DPS (chung2023diffusion) on celebA denoising. (a) NFE vs. PSNR plot, (b) step size (used in DPS only) vs. PSNR plot, (c) representative results by varying the NFE.
It is worth mentioning that Noise2Score (kim2021noise2score) is applicable for one-step denoising of the measurements by leveraging the Tweedie’s formula. However, as discussed in Sec. 6, APS is generally applicable to a wide class of inverse problems, whereas the applicability of Noise2Score is limited.
A.4.2 Generalizability across datasets and Blind Inverse Problems
Figure 6: Figure demonstrating the visual results of robustness and generalization ability of APS. First and second rows show the results on unseen validation data, and third row depicts generalization to another dataset. Corresponding quantitative analysis can be found in Tab. 5.
[…]
Table 5: APS is robust against unseen data samples and even generalizable accross different datasets once it is optimized. For the first $2$ columns, we use validation part of datasets and feed-forward our CNF on this totally unseen data. For the third column, we even show that pre-optimized APS can be leveraged to restore back the noised data samples from across various datasets. All the quantitative results align with their counterparts in Tab. 2 that confirms robustness and generalization ability of our pipeline which was not possible before.
We further observe that our optimized framework can be used on unseen data as well. Tab. 5 and Fig. 6 depicts that we achieve similar quantitative and qualitative reconstruction results when we sample from unseen celebA or ERA5 validation datasets. Note that score network is trained on train signals and CNF optimization has been conducted on test signals, i.e. validation is totally hidden to both teacher and student models. More strongly, our approach can also be leveraged on various datasets as pre-optimized inverse solver. To this end, we use our celebA optimized CNF model to perform Denoising task on the FFHQ (karras2019style) dataset. Same Table and Figure show that our model can remove the noise artifacts with a similar performance as it does on original data, confirming the generalizibility feature.
We further observed that APS can work in the absence of forward operator. In other words, we can perform blind inverse problems through our amortized posterior sampling. We used various imputation levels between 30% to 60% for ERA5 dataset, and conducted experiments with random choice of imputation in a blind manner. As a result, Fig. 7 shows that results as good as the original inverse solver with the known forward operation (at least $33$ PSNR across all different blind imputation levels).